# Word-Level Emotional Expression Control in Zero-Shot Text-to-Speech Synthesis

### v1

@@ 4.2.2 Ablation Study
To examine the effectiveness of our dynamic speaking rate control, we remove this component from the 1st-stage model ("w/o speed control"). As shown in Table 4, DNSV slightly increases from 4.980 to 5.067, and performance drops are observed across most expressive metrics, such as AutoPCP (2.650 to 2.499) and Emo2v. (0.866 to 0.844). This suggests that speaking rate variation provides important prosodic cues for emotional expression in TTS.
@@ Conclusion
Appendix A Details of CosyVoice2
Appendix B Speed Control
Appendix C Generation of Emotionally Varying Texts
We employ GPT-4o [41] to generate the corpus of sentences containing intra-sentence emotional transitions. To ensure the emotional transitions are contextually plausible, we construct prompts based on predefined scenarios, character relationships, and conversation topics. An example of the prompt is shown in Listing 1. Specifically, we first create a large pool of randomly generated environments, contexts, and interpersonal relationships with personality traits. During generation, three elements are randomly selected and injected into the prompt to guide GPT-4o in producing scripts.
Appendix D Data Filtering in Self-Training
Appendix E Predefined Emotion Attention Bias
Appendix F Evaluation Setup
F.1 Details of Dataset
We use the train-set, dev-set, and test-set of ESD666https://github.com/HLTSingapore/Emotional-Speech-Data [21] for training and evaluation. This dataset contains 350 parallel utterances, averaging 2.9 seconds in duration, spoken by 20 speakers: 10 native English and 10 native Mandarin (5 male and 5 female for each language). Each speaker expresses five emotions: happy, sad, neutral, angry, and surprised. All audio is sampled at 16 kHz and stored as 16-bit files.
For evaluation, we generate 1,000 emotion-speed-varying text samples (500 in Chinese and 500 in English) using the script provided in Appendix C. For each text sample, we randomly select emotional prompts from the ESD test set to match the emotion transitions required by the sentence. All emotion prompts within a single sentence are drawn from the same speaker to ensure consistency. The reference audio for the target speaker is also randomly selected from the same language-speaker subset. As a result, approximately 1 out of every 5 samples features emotional prompts and a target speaker from the same speaker-emotion setting, given that the ESD dataset contains 5 emotions.
F.2 Baselines
We adopt four strong zero-shot TTS systems as baselines:
F.3 Evaluation Metrics
Objective Metrics
Subjective Evaluation
Appendix G Out-of-Domain Evaluation
To evaluate the generalization ability of our approach under an out-of-domain dataset, we conduct word-level emotion and speaking rate control experiments on the CASIA dataset [68], as organized according to Appendix F. The CASIA corpus is a Mandarin emotional speech dataset recorded by four native speakers and covers six emotion categories: neutral, angry, fear, happy, sad, and surprise. Some of these emotions are not seen during training, which makes CASIA suitable for testing the cross-domain robustness of controllable speech synthesis. The results are shown in Table 5. Our method, WeSCon, demonstrates strong performance across nearly all evaluation metrics, achieving lower DNSV and higher speaker similarity (S-SIM), emotional similarity (Emo2vec), and arousal scores compared to other baselines. The overall results are consistent with those in Table 1, further confirming that our method generalizes well to unseen speakers and novel emotional patterns.
| Method |
DNSV | S-SIM |
Emotion | |||
| Emo2v. | Aro. | |||||||
| Index-TTS | 1.217 | 8.887 | 0.468 | 2.444 | 0.824 | 0.502 | ||
| F5-TTS | 1.374 | 8.940 | 0.462 | 2.539 | 0.845 | 0.526 | ||
| Spark-TTS | 1.299 | 8.720 | 0.439 | 2.496 | 0.841 | 0.523 | ||
| CosyVoice2 | 1.405 | 8.093 | 0.542 | 2.503 | 0.835 | 0.517 | ||
| WeSCon (1st) | 1.411 | 4.680 | 0.587 | 2.670 | 0.869 | 0.548 | ||
| WeSCon (2nd) | 1.478 | 4.641 | 0.590 | 2.624 | 0.867 | 0.552 | ||
Appendix H Training Progress
@@
Justification: We conducted crowdsourced experiments, and Appendix F.3 includes the full text of the instructions provided to participants, along with illustrative screenshots. No compensation was provided to the annotators.
@@
Justification: We use a large language model (LLM) to augment our experimental dataset. The specific usage of the LLM is described in detail in Appendix C. Although the LLM is not a core model component, its use contributes directly to the experimental design and data quality, and thus we provide a clear explanation of its role.

### v2

@@ 4.2.2 Ablation Study
To examine the effectiveness of our dynamic speaking rate control, we remove this component from the 1st-stage model ("w/o speed control"). As shown in Table 4, DNSV slightly increases from 4.980 to 5.067, and performance drops are observed across most expressive metrics, such as AutoPCP (2.650 to 2.499) and Emo2v. (0.866 to 0.844). This suggests that speaking rate variation provides important prosodic cues for emotional expression in TTS. In addition, we further investigate the interaction between speaking rate control and emotional expression in Appendix C.
@@ Conclusion
6 Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant U23B2053 and Grant 62176182 and the Open Research Fund from Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ) (No.GML-KF-24-16).
@@ 6 Acknowledgements
Justification: We conducted crowdsourced experiments, and Appendix 9 includes the full text of the instructions provided to participants, along with illustrative screenshots. No compensation was provided to the annotators.
@@
Justification: We use a large language model (LLM) to augment our experimental dataset. The specific usage of the LLM is described in detail in Appendix D. Although the LLM is not a core model component, its use contributes directly to the experimental design and data quality, and thus we provide a clear explanation of its role.
@@
Appendix A Details of CosyVoice2
Appendix B Speed Control
Appendix C Interaction Between Speaking Rate and Emotion
This section further examines the relationship between speaking rate and emotional expression. Speaking rate and emotional state are strongly coupled in human speech, as different emotions are typically associated with distinct prosodic rhythms and energy patterns. Since both rate and emotional cues are derived from the same prompt speech, temporal resampling inevitably alters the perceived emotional expression. To quantify the effect of speaking rate on perceived emotion, 100 emotional utterances are randomly selected from the test set as prompts. The resampling ratio is systematically varied from 0.5 to 2.0 in increments of 0.25, using the same target text for all conditions, as summarized in Table 5. For each condition, the emotion similarity between the generated speech and both the original and rate-matched (re-rated) reference speech is calculated using the Emo2v. score.
| Resampling Ratio | Emo2v. | Emo2v. (Re-rated) |
| 0.5 (downsampled to half, speed up) | 0.57 | 0.86 |
| 0.75 | 0.85 | 0.88 |
| 1.0 | 0.90 | 0.90 |
| 1.25 | 0.83 | 0.89 |
| 1.5 | 0.75 | 0.90 |
| 1.75 | 0.68 | 0.91 |
| 2.0 (interpolated to twice, speed down) | 0.51 | 0.87 |
The results show that emotion similarity declines substantially when the reference is not rate-matched, whereas it remains stable when compared with rate-adjusted references. Notably, when the resampling ratio deviates significantly from the natural range (e.g., 0.75 ~1.5), the perceived emotion becomes less consistent, likely due to distortion of spectral dynamics and pitch contours caused by excessive time-stretching or compression. These findings confirm that speaking rate provides essential prosodic cues for emotional perception, consistent with the observations in Section 4.2.2.
Appendix D Generation of Emotionally Varying Texts
We employ GPT-4o [28] to generate the corpus of sentences containing intra-sentence emotional transitions. To ensure the emotional transitions are contextually plausible, we construct prompts based on predefined scenarios, character relationships, and conversation topics. An example of the prompt is shown in Listing LABEL:prompt. Specifically, we first create a large pool of randomly generated environments, contexts, and interpersonal relationships with personality traits. During generation, three elements are randomly selected and injected into the prompt to guide GPT-4o in producing scripts.
Appendix E Data Filtering in Self-Training
Appendix F Predefined Emotion Attention Bias
Appendix G Evaluation Setup
G.1 Details of Dataset
We use the train-set, dev-set, and test-set of ESD666https://github.com/HLTSingapore/Emotional-Speech-Data [66] for training and evaluation. This dataset contains 350 parallel utterances, averaging 2.9 seconds in duration, spoken by 20 speakers: 10 native English and 10 native Mandarin (5 male and 5 female for each language). Each speaker expresses five emotions: happy, sad, neutral, angry, and surprised. All audio is sampled at 16 kHz.
For evaluation, we generate 1,000 emotion-speed-varying text samples (500 in Chinese and 500 in English) using the script provided in Appendix D. For each text sample, we randomly select emotional prompts from the ESD test set to match the emotion transitions required by the sentence. All emotion prompts within a single sentence are drawn from the same speaker to ensure consistency. The reference audio for the target speaker is also randomly selected from the same language-speaker subset. As a result, approximately 1 out of every 5 samples features emotional prompts and a target speaker from the same speaker-emotion setting, given that the ESD dataset contains 5 emotions.
G.2 Baselines
We adopt four strong zero-shot TTS systems as baselines:
G.3 Evaluation Metrics
Objective Metrics
Subjective Evaluation
Appendix H Out-of-Domain Evaluation
To evaluate the generalization ability of our approach under an out-of-domain dataset, we conduct word-level emotion and speaking rate control experiments on the CASIA dataset [38], as organized according to Appendix G. The CASIA corpus is a Mandarin emotional speech dataset recorded by four native speakers and covers six emotion categories: neutral, angry, fear, happy, sad, and surprise. Some of these emotions are not seen during training, which makes CASIA suitable for testing the cross-domain robustness of controllable speech synthesis. The results are shown in Table 6. Our method, WeSCon, demonstrates strong performance across nearly all evaluation metrics, achieving lower DNSV and higher speaker similarity (S-SIM), emotional similarity (Emo2vec), and arousal scores compared to other baselines. The overall results are consistent with those in Table 1, further confirming that our method generalizes well to unseen speakers and novel emotional patterns.
| Method | WER/CER | DNSV | S-SIM | AutoPCP | Emotion | |
| Emo2v. | Aro. | |||||
| Index-TTS | 1.217 | 8.887 | 0.468 | 2.444 | 0.824 | 0.502 |
| F5-TTS | 1.374 | 8.940 | 0.462 | 2.539 | 0.845 | 0.526 |
| Spark-TTS | 1.299 | 8.720 | 0.439 | 2.496 | 0.841 | 0.523 |
| CosyVoice2 | 1.405 | 8.093 | 0.542 | 2.503 | 0.835 | 0.517 |
| WeSCon (1st) | 1.411 | 4.680 | 0.587 | 2.670 | 0.869 | 0.548 |
| WeSCon (2nd) | 1.478 | 4.641 | 0.590 | 2.624 | 0.867 | 0.552 |
Appendix I Training Progress